OpenStudy (anonymous):

If a triangle contains two 60° angles, then it is equilateral

OpenStudy (anonymous):

I say true

OpenStudy (rsadhvika):

yes third angle wd be 60 also becoz sum of angles in triangle is 180 60 + 60 + x = 180 x = 180 - 120 = 60

OpenStudy (rsadhvika):

since all angles are equal it is equilateral triangle

OpenStudy (unklerhaukus):